# The evolution of trust as a cognitive shortcut in repeated interactions

### v2

@@ 4 Discussion
Third, we have modelled a well-mixed population of agents that are all similar in their cognitive capabilities. It would be important to examine the effects of spatial structure [51], which can facilitate the evolution of cooperation in repeated games such as the Prisoner’s Dilemma [23]. Moreover, agents might differ in how easily they can verify another agent’s actions, and in the complexity of strategies that they can use. For example, when considering interactions between humans and AI agents, the model could be extended to allow AI agents to verify actions at lower costs, and to use more complex strategies such as modelling and predicting human-partner trust behaviours. This would represent AI agents’ greater access to information and processing power, respectively. This could be achieved by modelling a separate population of AI agents that compete with and copy strategies from each other, but earn payoffs from playing with a population of human agents.
Finally, the model of trust as a cognitive shortcut to avoid the cost of verification could apply beyond interactions that are repeated between the same pair of agents. These include $n$-player social dilemmas such as repeated public goods games. Trust-based strategies could also apply with indirect reciprocity, in cases where it is costly to get information from third parties about a partner’s previous actions.
In conclusion, we have developed a behavioural measure of trust applies across the spectrum of symmetric two-player social dilemma games, and that distinguishes observed trust from observed cooperation. Because this behavioural measure does not depend on access to an agent’s internal state, it can apply to both human and AI agents. Using the methods of evolutionary game theory, we have shown that Trust-based Cooperation, which verifies its partners’ actions only occasionally after a trust-threshold is reached, is a successful strategy in a wide range of social interactions. We have also shown that the presence of trust increases the overall level of cooperation in a population, especially when agents can make unintentional errors in their actions. This adds formal support for the idea that high-trust societies lead to more cooperative and friction-free economies [1].
@@
The authors used an AI-based tool (ChatGPT 5.1 with Deep Research assistance) solely to support the identification of relevant literature on costly monitoring in classic game theory and behavioural economics. All cited sources were independently reviewed and integrated by the authors. No AI-generated text was used in the submitted manuscript.

### v3

@@ 4 Discussion
Third, we have modelled a well-mixed population of agents that are all similar in their cognitive capabilities. It would be important to examine the effects of spatial structure [58], which can facilitate the evolution of cooperation in repeated games such as the Prisoner’s Dilemma [31]. Moreover, agents might differ in how easily they can monitor another agent’s actions, and in the complexity of strategies that they can use. For example, when considering interactions between humans and AI agents, the model could be extended to allow AI agents to monitor actions at lower costs, and to use more complex strategies such as modelling and predicting human-partner trust behaviours. This would represent AI agents’ greater access to information and processing power, respectively. This could be achieved by modelling a separate population of AI agents that compete with and copy strategies from each other, but earn payoffs from playing with a population of human agents.
Finally, the game-theoretic formalisation of trust as a cognitive shortcut to avoid the cost of continuous monitoring could apply beyond interactions that are repeated between the same pair of agents. These include $n$-player social dilemmas such as repeated public goods games. Trust-based strategies could also apply with indirect reciprocity, in cases where it is costly to get information from third parties about a partner’s previous actions.
In conclusion, we have developed a behavioural measure of trust that applies across the spectrum of symmetric two-player social dilemma games, and that distinguishes observed trust from observed cooperation. Because this behavioural measure does not depend on access to an agent’s internal state, it can apply to both human and AI agents. Using the methods of evolutionary game theory, we have shown that Trust-based Cooperation, which verifies its partners’ actions only occasionally after a trust-threshold is reached, is a successful strategy in a wide range of social interactions. We have also shown that the presence of trust increases the overall level of cooperation in a population, especially when agents can make unintentional errors in their actions. This adds formal support for the idea that high-trust societies lead to more cooperative and friction-free economies [1].
@@
The authors used AI-based tools (ChatGPT 5.1 and Elicit) solely to assist with the identification of potentially relevant literature. All cited sources were independently reviewed and evaluated by the authors. AI tools were not used for data analysis, interpretation, or generation of manuscript text.